OpenStudy (anonymous):

Using elimination: 2x + 5y = -5 x + 3y = 3

OpenStudy (anonymous):

2x + 5y = -5 x + 3y = 3 *2 thn 2x + 5y = -5 2x + 3y = 3 - - - ------------- 0x+2y=-8 y=-4 and put y is in 2x + 5y = -5 and get the reslt